the demand and supply equations for a certain item are given by
D= -5p+40
s= -p+30p-8
Find the equilibrium price.

Answer by elima(1433)

the demand and supply equations for a certain item are given by
D= -5p+40
s= -p+30p-8
Find the equilibrium price.
when they equal each other they are in equalibrium
-5p+40=-p+30p-8
40+8=-p+30p+5p
48=34p
$1.41
:)